# DIAGNOSIS: Detecting unauthorized data usages in text-to-image diffusion models

### 04 Feb 2024 v2

@@ 1 Introduction
Recently, text-to-image diffusion models have showcased outstanding capabilities in generating a wide range of high-quality images. Notably, the release of Stable Diffusion (Rombach et al., 2022), one of the most advanced and open-source text-to-image-diffusion models, has significantly contributed to this progress. There has been a remarkable surge in the applications that use Stable Diffusion as the foundation model. Consequently, more users adopt these models as tools for generating images with rich semantics according to their preferences. As diffusion models become prevalent, the problems related to the responsible development of them become more and more  (Wen et al., 2023; Fernandez et al., 2023; Wang et al., 2023a; Cui et al., 2023; Zhao et al., 2023).
@@
Inspired by recent findings that the text-to-image diffusion models can memorize contents in the training data (Carlini et al., 2023a; Somepalli et al., 2023a; b), in this paper, we solve the unauthorized data usages detection problem from the lens of the memorizations. Existing works about the memorization of the text-to-image diffusion models focus on the memorization of whole samples (sample-level memorization (Carlini et al., 2023a; Somepalli et al., 2023a; b)). A model is considered to have sample-wise memorization if a specific sample can be accurately identified as a training sample of the model through membership inference attacks (Carlini et al., 2023a). Thus, an intuitive way is exploiting membership inference techniques (Shokri et al., 2017; Chen et al., 2020) to detect if specific data are used to train or fine-tune the given model. However, Duan et al. (2023) demonstrate that existing membership inference methods are ineffective for text-to-image diffusion models such as Stable Diffusion. For example, the state-of-the-art membership inference method for the diffusion model (i.e., SecMI (Duan et al., 2023)) only achieves 66.1% success rate for the membership inference on the stable-diffusion-v1-5 model (Rombach et al., 2022) under white-box setting. Performing membership inference for large diffusion models in practical black-box settings is even more challenging (Dubiński et al., 2023). Different from the sample-level memorization, in this work, we focus on diffusion models’ memorization on specific elements in the training data and propose an approach for detecting unauthorized data usages via planting the injected element-level memorizations into the model trained or fine-tuned on the protected dataset by modifying the protected training data. More specifically, when a set of protected images are uploaded to the internet, it will be processed by a specific function (called signal function) that is stealthy to humans but can be captured and memorized for diffusion models. Therefore, after the models are trained or fine-tuned on the “coated images” (i.e., images processed by the signal function), it will memorize the added signal function so that the unauthorized data usages can be detected by using a binary classifier (called signal classifier) to analyze if the given model has the memorization on the signal function (i.e., if the images generated by the model contains the signal function). Our method is independent of the model used in the unauthorized training or fine-tuning process, and it only has a small influence on the authorized training. To the best of our knowledge, this is the first work to study the more fine-grained element-level memorization for text-to-image diffusion models. Based on our design, we implemented our prototype called DIAGNOSIS (Detecting unauthorIzed dAta usaGes iN text-tO-image diffuSIon modelS) in PyTorch, experiments on mainstream text-to-image diffusion models (i.e., Stable Diffusion v1, Stable Diffusion v2 (Rombach et al., 2022), VQ Diffusion (Gu et al., 2022)) and popular model training or fine-tuning methods (i.e., Low-Rank Adaptation (LoRA) (Hu et al., 2022), DreamBooth (Ruiz et al., 2023), and standard training) demonstrate that our method is highly effective. It achieves 100.0% detection accuracy under various settings. Meanwhile, our proposed method has small influence on the generation quality of the models, and the “coated images” in the protected dataset are close to the original images. For example, Fig. 1 shows the visualizations of the generated samples by the standard models and the models planted with injected memorization. In Fig. 1, all images generated by the injected model are recognized as “contains signal function” by the signal classifier, but their distributional difference to the images generated by standard models is small. More visualizations of the generated samples can be found in Fig. 2, Fig. 3 and Fig. 4. The visualizations of the original training samples and their coated version can be found in Fig. 5.
@@ 3.2 Data Usages Tracing and Injected Memorization
Injected Memorization. Our idea is planting some unique behaviors into the models trained or fine-tuned on the protected dataset via modifying the dataset, then we can detect the unauthorized data usages by checking the behaviors of the inspected models. The research question we want to answer is “How to plant the unique behaviors into the models that trained or fine-tuned on the protected dataset?”. Recent research (Carlini et al., 2023a; Somepalli et al., 2023a; b) find that the text-to-image diffusion models have strong memorizations on the duplicated training samples. That is, the model tends to produce the images that are highly similar to some memorized training samples. Such behaviors are dependent on the training data, and they are highly orthogonal to the learning algorithms. Inspired by these observations, we propose to plant the unique behaviors into the models by modifying the protected dataset and injecting extra memorizations on some unique contents such as a stealthy image processing function (i.e., signal function). We call such memorizations as the injected memorization. Different from existing studies (Carlini et al., 2023a; Somepalli et al., 2023a; b) that focus on the sample-level memorization, we focus on the more fine-grained element-level memorization of specific elements (e.g., an image processing function) for text-to-image diffusion models. In this paper, we denote signal function $\mathcal{S}$ as the process to add the selected unique content into the image. We use $\mathcal{O}_{\mathcal{S}}$ to denote the set of image samples processed by signal function $\mathcal{S}$ and call $\mathcal{O}_{\mathcal{S}}$ as the signal images. The injected memorizations can be conditioned on a text trigger function $\eta$ in the text prompt. That means the model tends to generate signal images if the function $\eta$ is applied to the text prompt, and the model does not have such preference otherwise. It can also be unconditional. Formally, we define injected memorization as follows:
@@ 4.1 Experiment Setup
Implementation Details. By default, the coating rate we used for unconditional injected memorization and the trigger-conditioned injected memorization are 100.0% and 20.0%, respectively. We use 50 text prompts to approximate the memorization strength (i.e., $N=50$) by default. The default warping strength are 2.0 and 1.0 for unconditional injected memorization and trigger-conditioned injected memorization, respectively. The default hyper-parameters for the hypothesis testing are discussed in § 3.3. The trigger-conditioned injected memorization is corresponded to the Scenario I described in § 3.1, and the unconditional injected memorization is corresponded to both Scenario I and Scenario II. Our code will be released upon publication.
@@ 4.3 Ablation Study
Different Warping Strengths. As we discussed in § 3.2, we use the image warping function as the signal function of the injected memorization. The effects of the image warping function is
@@
controlled by the hyper-parameter warping strength, which is defined as the scale of the warping-induced perturbations (i.e., $s$ in Eq. 5 in § A.1). We investigate the influence of different warping strengths for image warping function in the coating process.
@@ 4.4 Discussion about different types of injected memorization
5 Ethic Statement
The study of machine learning’s security and privacy aspects has the potential to give rise to ethical concerns (Carlini et al., 2023b; Wang et al., 2023b; Thudi et al., 2022; Liu et al., 2022). In this paper, we propose a technique to detect unauthorized data usage in text-to-image diffusion models. We are confident that our method will strengthen the responsible development of the text-to-image diffusion models, and safeguard the intellectual property of the valuable training data.
@@ A.9 Results on Textual Inversion
Table 14: Effectiveness on Textual Inversion (Gal et al., 2023)
@@ A.11 Symbol Table
Table 15: Summary of Symbols

### 09 Apr 2024 v3

@@ 1 Introduction
Recently, text-to-image diffusion models have showcased outstanding capabilities in generating a wide range of high-quality images. Notably, the release of Stable Diffusion (Rombach et al., 2022), one of the most advanced and open-source text-to-image-diffusion models, has significantly contributed to this progress. There has been a remarkable surge in the applications that use Stable Diffusion as the foundation model. Consequently, more users adopt these models as tools for generating images with rich semantics according to their preferences. As diffusion models become prevalent, the problems related to the responsible development of them become increasingly critical  (Wen et al., 2023; Fernandez et al., 2023; Wang et al., 2023a; Cui et al., 2023; Zhao et al., 2023).
@@
Inspired by recent findings that the text-to-image diffusion models can memorize contents in the training data (Carlini et al., 2023a; Somepalli et al., 2023a; b; Wen et al., 2024), in this paper, we solve the unauthorized data usages detection problem from the lens of the memorizations. Existing works about the memorization of the text-to-image diffusion models focus on the memorization of whole samples (sample-level memorization (Carlini et al., 2023a; Somepalli et al., 2023a; b)). A model is considered to have sample-wise memorization if a specific sample can be accurately identified as a training sample of the model through membership inference attacks (Carlini et al., 2023a). Thus, an intuitive way is exploiting membership inference techniques (Shokri et al., 2017; Chen et al., 2020) to detect if specific data are used to train or fine-tune the given model. However, Duan et al. (2023) demonstrate that existing membership inference methods are ineffective for text-to-image diffusion models such as Stable Diffusion. For example, the state-of-the-art membership inference method for the diffusion model (i.e., SecMI (Duan et al., 2023)) only achieves 66.1% success rate for the membership inference on the stable-diffusion-v1-5 model (Rombach et al., 2022) under white-box setting. Performing membership inference for large diffusion models in practical black-box settings is even more challenging (Dubiński et al., 2023). Different from the sample-level memorization, in this work, we focus on diffusion models’ memorization on specific elements in the training data and propose an approach for detecting unauthorized data usages via planting the injected element-level memorizations into the model trained or fine-tuned on the protected dataset by modifying the protected training data. More specifically, when a set of protected images are uploaded to the internet, it will be processed by a specific function (called signal function) that is stealthy to humans but can be captured and memorized for diffusion models. Therefore, after the models are trained or fine-tuned on the “coated images” (i.e., images processed by the signal function), it will memorize the added signal function so that the unauthorized data usages can be detected by using a binary classifier (called signal classifier) to analyze if the given model has the memorization on the signal function (i.e., if the images generated by the model contains the signal function). Our method is independent of the model used in the unauthorized training or fine-tuning process, and it only has a small influence on the authorized training. To the best of our knowledge, this is the first work to study the more fine-grained element-level memorization for text-to-image diffusion models. Based on our design, we implemented our prototype called DIAGNOSIS (Detecting unauthorIzed dAta usaGes iN text-tO-image diffuSIon modelS) in PyTorch, experiments on mainstream text-to-image diffusion models (i.e., Stable Diffusion v1, Stable Diffusion v2 (Rombach et al., 2022), VQ Diffusion (Gu et al., 2022)) and popular model training or fine-tuning methods (i.e., Low-Rank Adaptation (LoRA) (Hu et al., 2022), DreamBooth (Ruiz et al., 2023), and standard training) demonstrate that our method is highly effective. It achieves 100.0% detection accuracy under various settings. Meanwhile, our proposed method has small influence on the generation quality of the models, and the “coated images” in the protected dataset are close to the original images. For example, Fig. 1 shows the visualizations of the generated samples by the standard models and the models planted with injected memorization. In Fig. 1, all images generated by the injected model are recognized as “contains signal function” by the signal classifier, but their distributional difference to the images generated by standard models is small. More visualizations of the generated samples can be found in Fig. 2, Fig. 3 and Fig. 4. The visualizations of the original training samples and their coated version can be found in Fig. 5.
@@ 3.2 Data Usages Tracing and Injected Memorization
Injected Memorization. Our idea is planting some unique behaviors into the models that were trained or fine-tuned on the protected dataset via modifying the dataset, then we can detect the unauthorized data usages by checking the behaviors of the inspected models. The research question we want to answer is “How to plant the unique behaviors into the models that were trained or fine-tuned on the protected dataset?”. Recent research (Carlini et al., 2023a; Somepalli et al., 2023a; b) find that the text-to-image diffusion models have strong memorizations on the duplicated training samples. That is, the model tends to produce the images that are highly similar to some memorized training samples. Such behaviors are dependent on the training data, and they are highly orthogonal to the learning algorithms. Inspired by these observations, we propose to plant the unique behaviors into the models by modifying the protected dataset and injecting extra memorizations on some unique contents such as a stealthy image processing function (i.e., signal function). We call such memorizations as the injected memorization. Different from existing studies (Carlini et al., 2023a; Somepalli et al., 2023a; b) that focus on the sample-level memorization, we focus on the more fine-grained element-level memorization of specific elements (e.g., an image processing function) for text-to-image diffusion models. In this paper, we denote signal function $\mathcal{S}$ as the process to add the selected unique content into the image. We use $\mathcal{O}_{\mathcal{S}}$ to denote the set of image samples processed by signal function $\mathcal{S}$ and call $\mathcal{O}_{\mathcal{S}}$ as the signal images. The injected memorizations can be conditioned on a text trigger function $\eta$ in the text prompt. That means the model tends to generate signal images if the function $\eta$ is applied to the text prompt, and the model does not have such preference otherwise. It can also be unconditional. Formally, we define injected memorization as follows:
@@ 4.1 Experiment Setup
Implementation Details. By default, the coating rate we used for unconditional injected memorization and the trigger-conditioned injected memorization are 100.0% and 20.0%, respectively. We use 50 text prompts to approximate the memorization strength (i.e., $N=50$) by default. The default warping strength are 2.0 and 1.0 for unconditional injected memorization and trigger-conditioned injected memorization, respectively. The default hyper-parameters for the hypothesis testing are discussed in § 3.3. The trigger-conditioned injected memorization is corresponding to the Scenario I described in § 3.1, and the unconditional injected memorization is corresponding to both Scenario I and Scenario II. Our code will be released upon publication.
@@ 4.3 Ablation Study
Different Warping Strengths. As we discussed in § 3.2, we use the image warping function as the
@@
signal function of the injected memorization. The effects of the image warping function is controlled by the hyper-parameter warping strength, which is defined as the scale of the warping-induced perturbations (i.e., $s$ in Eq. 5 in § A.1). We investigate the influence of different warping strengths for image warping function in the coating process.
@@ 5 Conclusion
Ethic Statement
The study of machine learning’s security and privacy aspects has the potential to give rise to ethical concerns (Carlini et al., 2023b; Wang et al., 2023b; Thudi et al., 2022; Tao et al., 2023; Liu et al., 2022; Tao et al., 2022). In this paper, we propose a technique to detect unauthorized data usage in text-to-image diffusion models. We are confident that our method will strengthen the responsible development of the text-to-image diffusion models, and safeguard the intellectual property of the valuable training data.
Acknowledgement
We thank the anonymous reviewers for their valuable comments. This research is supported by Sony AI, IARPA TrojAI W911NF-19-S-0012, NSF 2342250 and 2319944. It is also partially funded by research grants to D. Metaxas through NSF 2310966, 2235405, 2212301, 2003874, and AFOSR-835531. Any opinions, findings, and conclusions expressed in this paper are those of the authors only and do not necessarily reflect the views of any funding agencies.
@@ A.9 Results on Textual Inversion
Table 14: Effectiveness on Textual Inversion (Gal et al., 2023).
@@ A.11 Symbol Table
Table 15: Summary of symbols.